In what direction a line be drawn through the point (1,2) so that its point of intersection with the line x+y=4 is at a distance


(root 6/3) from the given point.


A. 75 degree                              B.60 degree


C. 45 degree                              D.30 degree

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Anindita

let parametric form of line

x=1+ √6/3 cosθ

y=2+√6/3 sin θ

if this line cut the given line

then

 1+ √6/3 cosθ  +  2+√6/3 sin θ =4

 √6/3 (cosθ  +  sin θ)=1

 sin (θ+45) =Sin60

 so  θ+45 =60   or 120

or    θ  =15 or  75


 Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE & AIEEE preparation. 

 All the best. 
 
Regards, 
Askiitians Experts 
Badiuddin